# Learning Mutual Excitation for Hand-to-Hand and Human-to-Human Interaction Recognition

### v1

@@ Abstract
Recognizing interactive actions, including hand-to-hand interaction and human-to-human interaction, has attracted increasing attention for various applications in the field of video analysis and human-robot interaction. Considering the success of graph convolution in modeling topology-aware features from skeleton data, recent methods commonly operate graph convolution on separate entities and use late fusion for interactive action recognition, which can barely model the mutual semantic relationships between pairwise entities. To this end, we propose a mutual excitation graph convolutional network (me-GCN) by stacking mutual excitation graph convolution (me-GC) layers. Specifically, me-GC uses a mutual topology excitation module to firstly extract adjacency matrices from individual entities and then adaptively model the mutual constraints between them. Moreover, me-GC extends the above idea and further uses a mutual feature excitation module to extract and merge deep features from pairwise entities. Compared with graph convolution, our proposed me-GC gradually learns mutual information in each layer and each stage of graph convolution operations. Extensive experiments on a challenging hand-to-hand interaction dataset, i.e., the Assembely101 dataset, and two large-scale human-to-human interaction datasets, i.e., NTU60-Interaction and NTU120-Interaction consistently verify the superiority of our proposed method, which outperforms the state-of-the-art GCN-based and Transformer-based methods.
@@ I Introduction
Interactive action recognition is an important task, as it is an indispensable stream of action recognition [1, 2, 3, 4, 5] that has applications in video analysis [6, 7, 8, 9, 10, 11] and human-robot interaction [12, 13, 14, 15]. Recently, skeleton-based interactive action recognition has attracted increasing attention for the development of depth sensors, e.g., Kinect [16] and Intel RealSense [17], and the robustness of skeleton data to complicated backgrounds and viewpoint variations [18, 19]. We focus on interaction between pairwise entities, including hand-to-hand and human-to-human interaction.
@@ IV-A Datasets and Settings
Settings. All experiments are conducted on one Tesla V100 GPU with PyTorch. We adopt the SGD algorithm [49] with a Nesterov momentum of 0.9 and weight decay of 0.0004 as the optimizer. To make the training procedure more stable, the first 5 epochs are used for warm-up. The initial learning rate is set to 0.01 and is divided by 10 at the 35th and 55th epochs. The training process is terminated at the 65th epoch, and the batch size is set to 32 for both datasets. As a prepossessing step, the global movements of each skeleton sequence are removed by subtracting skeleton joint coordinates from the 2nd skeleton joint coordinates from the first frame. Skeleton sequences with different temporal frames are resized to 64 frames. We set $r$ to 8 as default value for all datasets.
@@ IV-D Visualization
V Conclusion
We investigate the skeleton-based interactive action recognition task and presents a mutual excitation graph convolutional network (me-GCN) by stacking mutual excitation graph convolution (me-GC) layers. Specifically, our proposed me-GC layer adopts a mutual topology excitation (MTE) module for topology modeling and a mutual feature excitation (MFE) module for feature extraction. Beyond the standard graph convolution layer, our me-GC layer can learn mutual semantic relationships between interactive body parts, thus facilitating the encoding of critical interactive motion patterns. Compared with GCN and Transformer-based methods, we achieve state-of-the-art performances on two challenging human-to-human interaction datasets and one large-scale hand-to-hand interaction dataset.

### v2

@@ Abstract
Recognizing interactive actions, including hand-to-hand interaction and human-to-human interaction, has attracted increasing attention for various applications in the field of video analysis and human-robot interaction. Considering the success of graph convolution in modeling topology-aware features from skeleton data, recent methods commonly operate graph convolution on separate entities and use late fusion for interactive action recognition, which can barely model the mutual semantic relationships between pairwise entities. To this end, we propose a mutual excitation graph convolutional network (me-GCN) by stacking mutual excitation graph convolution (me-GC) layers. Specifically, me-GC uses a mutual topology excitation module to firstly extract adjacency matrices from individual entities and then adaptively model the mutual constraints between them. Moreover, me-GC extends the above idea and further uses a mutual feature excitation module to extract and merge deep features from pairwise entities. Compared with graph convolution, our proposed me-GC gradually learns mutual information in each layer and each stage of graph convolution operations. Extensive experiments on a challenging hand-to-hand interaction dataset, i.e., the Assembely101 dataset, and two large-scale human-to-human interaction datasets, i.e., NTU60-Interaction and NTU120-Interaction consistently verify the superiority of our proposed method, which outperforms the state-of-the-art GCN-based and Transformer-based methods. Code is provided in https://github.com/nkliuyifang/me-GCN.
@@ I Introduction
Interactive action recognition is an important task, as it is an indispensable stream of action recognition [1, 2, 3, 4, 5] that has applications in video analysis [6, 7, 8, 9, 10, 11, 12, 13] and human-robot interaction [14, 15, 16, 17]. Recently, skeleton-based interactive action recognition has attracted increasing attention for the development of depth sensors, e.g., Kinect [18] and Intel RealSense [19], and the robustness of skeleton data to complicated backgrounds and viewpoint variations [20, 21]. We focus on interaction between pairwise entities, including hand-to-hand and human-to-human interaction.
@@ IV-A Datasets and Settings
Settings. All experiments are conducted on one Tesla V100 GPU with PyTorch. We adopt the SGD algorithm [51] with a Nesterov momentum of 0.9 and weight decay of 0.0004 as the optimizer. To make the training procedure more stable, the first 5 epochs are used for warm-up. The initial learning rate is set to 0.01 and is divided by 10 at the 35th and 55th epochs. The training process is terminated at the 65th epoch, and the batch size is set to 32 for both datasets. To ensure fair comparisons, we adopt the same preprocessing steps as in previous methods. For the NTU60-Interaction and NTU120-Interaction datasets, we follow the procedures outlined in [52, 53], removing the global movements from each skeleton sequence by subtracting the coordinates of each skeleton joint from the coordinates of the second joint in the first frame. For the Assemble101 dataset, we use the original data provided by [47] without applying any additional preprocessing steps. Skeleton sequences with different temporal frames are resized to 64 frames. We set $r$ to 8 as default value for all datasets.
@@ IV-D Visualization
V Conclusion and Future Work
We investigate the skeleton-based interactive action recognition task and presents a mutual excitation graph convolutional network (me-GCN) by stacking mutual excitation graph convolution (me-GC) layers. Specifically, our proposed me-GC layer adopts a mutual topology excitation (MTE) module for topology modeling and a mutual feature excitation (MFE) module for feature extraction. Beyond the standard graph convolution layer, our me-GC layer can learn mutual semantic relationships between interactive body parts, thus facilitating the encoding of critical interactive motion patterns. Compared with GCN and Transformer-based methods, we achieve state-of-the-art performances on two challenging human-to-human interaction datasets and one large-scale hand-to-hand interaction dataset. Future work focus on combining our method with recent Transformer-based methods like [58].